Starting from rest, a 10.06-cm-diameter compact disk takes 8.5 s to reach its operating angular velocity of 2381 rpm. Assume that the angular acceleration is constant. The disk�s moment of inertia is 1.00×10-5 kg·m2.

How much torque is applied to the disk?
0.0003 N*m

How many revolutions does it make before reaching full speed?
Answer #1

Using rotational kinematics

wf = wo + at

249.34 = a*8.5

a = 29.33 rad/sec2

So,

d = 1/2 * a * t2

d = 1059.68 radians

so,

number of revolutions = 168.65 revolutions

number of revolutions = 169 revolutions or 168 revolutions